• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 12:28:09 on October 18, 2010

    Entry number 335469

    keyword=Owl shift summary

    Shift started taking elastic calibration data on 1st setting. Runs were 
    stopped frequently due to lost of sync. We finally managed to cumulated 
    up to 2h30 of, hopefully, synchronized data (including the ones taken 
    during previous swing).
    
    Around 3am, we tried to move to the 2nd elastic setting. Bogies didn't 
    work, GUI became unresponsive. Had to make an access to reboot the IOC.
    
    Took some data on elastic setting 2nd (mostly out of sync).
    
    Around 5.30, moved to elastic setting 3 (bogies worked fine this time). 
    Took data for the remaining of the shift.
    
    Beam was taken away at 6.40.
    
    LHRS magnets were ramped down, target moved to empty, and controlled 
    access requested.
    
    
    Run list
    ========
    - 1st setting: 7485, 7486, 7487, 7488, 7489, 7490, 7491, 7493, 7494, 
    7495, 7496
    - 2nd setting: 7497, 7498
    - 3rd setting: 7499